OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3). 9 8 11 6

mathslover (mathslover):

C 11

mathslover (mathslover):

\[\color{dark blue}{3^2+2=9+2=11}\]

OpenStudy (unklerhaukus):

\[g(x) = x^2 + 2\]\[g(3) =(3)^2+2=11 \]